OpenStudy (anonymous):

how do you put x+2y=8 into slope intercept form?

OpenStudy (anonymous):

solve for y ...

OpenStudy (anonymous):

example?

OpenStudy (anonymous):

@mathg8 can you show me an example?

OpenStudy (anonymous):

messed up ... sorry

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

x+2y = 10 2y = -x +10 y= -x/2+5

OpenStudy (anonymous):

y=mx+b ( slope m. y intercept b)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

.